What is the required asset turnover for a firm with a 12% profit margin, 50% equity, and a 40% dividend payout that wishes to grow at 6% using only internal funding?
Answer #1

Asset turnover = Sales/ Assets

Given sustainable growth rate, g = 6%

6% = ROE * retention ratio

6% = ROE * (1- 40%)

Thus, ROE = 6%/(1-40%) = 6%/ 60% = 10%

ROE = 10%

Using Dupont formula,

ROE = Asset turn over * Profit margin * Leverage

Leverage = Asset/ Equity

As Debt = 50%, Equity = 50%

Asset = 50% + 50% = 1 ( since debt = 50% and equity = 50%)

Asset / Equity = (1/0.5) = 2 (since debt = equity = 0.5* Asset or half of assets)

10% = Asset turnover * 12% * 2

Asset turnover = 0.4166 = 0.42